OpenStudy (anonymous):

if one cup is equal to 8 fluid ounces, how many fluid ounces are in 3 cups

OpenStudy (freckles):

\[\frac{1 \text{ cup }}{8 \text{ fluid ounces }}=\frac{3 \text{ cup}}{x \text{ fluid ounces }}\]

OpenStudy (freckles):

solve the equaqtoin 1/8=3/x

OpenStudy (anonymous):

??

OpenStudy (freckles):

equation

OpenStudy (anonymous):

oh ok

OpenStudy (freckles):

were the question marks there because of my spelling of equation sorry it was a type-o

OpenStudy (anonymous):

is it 512

OpenStudy (freckles):

\[\frac{1}{8}=\frac{3}{x} \text{ is the same as } 1(x)=3(8)\]

OpenStudy (freckles):

3(8) shouldn't be 512

OpenStudy (anonymous):

now im confused

OpenStudy (mathstudent55):

1 cup = 8 fluid ounces 3 * 1 cups = ____ * 8 fluid ounces

OpenStudy (anonymous):

do i add or multiply

OpenStudy (freckles):

3(8) means multiply 3+8 means add

OpenStudy (freckles):

we have 3(8)

OpenStudy (freckles):

not 3+8

OpenStudy (anonymous):

24

OpenStudy (freckles):

yes 3(8) is 24 not 512 good job

OpenStudy (freckles):

\[\frac{1}{8}=\frac{3}{x} \text{ is the same as } 1(x)=3(8) \\ \text{ so we have } x=3(8)=24\]

OpenStudy (anonymous):

i feel dumb

OpenStudy (mathstudent55):

1 cup = 8 fluid ounces 3 * 1 cups = ____ * 8 fluid ounces 3 * 1 cups = 3 * 8 fluid ounces 3 cups = 24 fluid ounces